Use the following information to answer the question. Round your answers if necessary, to two decimal places.

Firm A can acquire firm B for $78,750 in cash or with stock worth $78,750 priced at its current price of $25 per share of stock. The synergy value of the deal is $15,000. Both firms are 100% equity financed.

Firm A: Number of Shares = 10,000 ; Price per Share = $25.00

Firm B: Number of Shares = 10,000 ; Price per Share = $10.00

What is the cost of acquisition when stock financing is used?
Answer #1

The shareholders of firm B will get (78750/25) 3150 Shares of Firm B

So their holding percentage post acquisition will be 3150/(10000+3150)= 3150/13150=.2395 or 23.95%

The value of the firm A (Before acquisition) = Number of shares * share price = 10000*25= $250000

The value of the firm A (Before acquisition) = Number of shares * share price = 10000*10= $100000

Synergy value of acquisition = $ 15000

so

the Value of firm after acquisition will be

= value of firm A + value of firm B + synergy

= $250000+$100000+15000

= $365000

Value of Acquisition will be ($365000*23.95%)-$100000 = $87417.5-100000= -12582.5